OpenStudy (anonymous):

At the movie theatre, child admission is $6.00 and adult admission is $9.20 . On Wednesday, four times as many adult tickets as child tickets were sold, for a total sales of $1455.20 . How many child tickets were sold that day?

OpenStudy (kropot72):

Let the number of child tickets be x. Then the number of adult tickets is 4x. We can write the following equation: 6x + (9.2 * 4x) = 1455.2 which can be simplified to: 6x + 36.8x = 1455.2 ............(1) Now you just need to solve equation (1) to find the value of x. Can you do that?

OpenStudy (kropot72):

The first step in solving equation (1) is to add the two terms on the left hand side. Can you do that and post your result?

OpenStudy (anonymous):

@kropot72 I'm working on it

OpenStudy (anonymous):

i got 236.4

OpenStudy (kropot72):

6x + 36.8x = ? x Just add the two coefficients of x.

OpenStudy (kropot72):

So what is 6 + 36.8 ?

OpenStudy (anonymous):

@kropot72 42.8

OpenStudy (kropot72):

Correct! So equation (1) simplifies to: 42.8x = 1455.2 ............(2) Now if you divide both sides of equation (2) by 42.8 you will have the value of x: \[x=\frac{1455.2}{42.8}=you\ can\ calculate\]

OpenStudy (anonymous):

34 right @kropot72

OpenStudy (kropot72):

Good work! 34 child tickets were sold.

OpenStudy (anonymous):

thanks :) @kropot72

OpenStudy (kropot72):

You're welcome :)